OpenStudy (anonymous):

Simplify: (2^n - 6^n)/(1-3^n)

OpenStudy (mathmath333):

\(\rm \color{black}{ \dfrac{2^n-6^n}{1-3^n}}\) \(\rm \color{black}{= \dfrac{2^n-2^n\times3^n}{1-3^n}}\) \(\rm \color{black}{ =\dfrac{2^n(1-3^n)}{1-3^n}}\) \(\rm \color{black}{ =2^n}\)